# Regularized Adjusted Plus-Minus Models for Evaluating and Scouting Football (Soccer) Players using Possession Sequences

### v1

@@ Abstract
This paper presents a novel framework for evaluating players in association football (soccer). Our method uses possession sequences, i.e. sequences of consecutive on-ball actions, for deriving estimates for player strengths. On the surface, the methodology is similar to classical adjusted plus-minus rating models using mainly regularized regression techniques. However, by analyzing possessions, our framework is able to distinguish on-ball and off-ball contributions of players to the game. From a methodological viewpoint, the framework explores four different penalization schemes, which exploit football-specific structures such as the grouping of players into position groups as well as into common strength groups. These four models lead to four ways to rate players by considering the respective estimate of each model corresponding to the player. The ratings are used to analyze the 2017/18 season of the Spanish La Liga. We compare similarities as well as particular use cases of each of the penalized models and provide guidance for practitioners when using the individual model specifications. Finally, we conclude our analysis by providing a domain-specific statistical evaluation framework, which highlights the potential of the penalized regression approaches for evaluating players.
@@ 1 Introduction
In football, the decision-making process is traditionally realized by a qualitative assessment of players by professionals such as coaches, managers, or scouts. However, lately, there have been many attempts and advances toward a more objective approach to quantifying the impact of players on the game. For example, traditional statistics such as goals or pass completion percentage, which have been used extensively to analyze the strength of players, have been replaced by metrics such as expected goals (xG, cf. e.g. Robberechts and Davis, 2020; Anzer and Bauer, 2021) or passing models based on statistical approaches (Szczepański and McHale, 2016; Håland et al., 2019, 2020). In general, these novel data-driven methods can be categorized into two distinct groups: bottom-up and top-down rating systems (cf. Hvattum and Gelade, 2021). The former assesses player performances by assigning values to each action performed and then aggregating them for each player over the course of a relevant period (i.e. a match or a season). Top-down ratings on the other hand evaluate players by breaking down the whole team’s performance and distributing credit onto the players involved. Typically, bottom-up approaches use very granular data on an event level in order to assign values to either specific actions such as shots (xG) or passes or more generally any type of action (cf. Decroos et al., 2019). While bottom-up ratings are therefore able to assess a player’s quality down to each particular action performed by the athlete, they however lack the ability to account for team strengths as well as off-ball actions by teammates. In contrast, top-down ratings such as plus-minus (PM) models (see Hvattum, 2019 for an overview) rely only on information gained from specific segments of a match, which in general is much less detailed than event stream data. These approaches are by design not able to account for specific actions taken by a player. They are, however, able to incorporate effects of teammates.
@@
In this work, we present a novel framework for evaluating player strength. On the surface, the methodology is similar to plus-minus (PM) models as we mainly use penalized regression models in order to infer a rating of players (cf. Kharrat et al., 2020; Hvattum, 2019). However, we extend the classical PM approach in two ways. First, we separate matches into possessions, which are sequences of consecutive on-ball actions. Thus, we consider much more granular segments of the game than the usual PM models, where a segment is defined as one where the same set of players is on the field. The segmentation in possessions leads to a data structure where it is possible to separate effects of being part of a possession, i.e. actually performing on-ball actions, from effects of only being on the pitch. From a domain-specific viewpoint, this is highly attractive as it is well known that off-ball actions have a substantial impact in soccer (cf. Spearman, 2018; Fernández et al., 2021). Using possessions data, it is also necessary to adjust the dependent variable that measures the outcome of a segment. In the classical PM approach the (normalized) goal difference in a segment is used. Since possessions are way shorter segments, goal difference is not appropriate as response variable. Instead, we focus on whether the possession ends in a goal or not as our outcome variable. Second, we provide a variety of penalization structures, which have not been considered in this context. Most of the classical plus-minus models use a ridge regression penalty, i.e. an $L_{2}$-type of penalty on the regression coefficient (cf. e.g. Kharrat et al., 2020). In a more innovative approach Pantuso and Hvattum (2021) adjust the penalty term by accounting for an age effect as well as a similarity effect within the shrinkage procedure. However, to the best of our knowledge, there has been no comparison or analysis of penalization structures beyond (adaptions of) the ridge penalty. We analyze several regularization structures that exploit the implicit grouping of players that arises from the different position groups on the field. In this way, our player evaluations are able to provide interesting and novel insights into player strengths where different penalties may be used for different use cases from practitioners. Finally, our last contribution aims at objectively evaluating ratings. In order to measure the relevancy of the novel modeling approaches presented in this work, we conduct a stastical evaluation specifically suited for football. The idea is similar to the validity experiments provided in Hvattum and Gelade (2021). That is, we measure whether a rating is superior to another rating by analyzing the ability of improving match results predictions using only the information from the rating. Besides comparing the ratings for the different regularized models, we also present comparisons to reference models using different kinds of data (ELO team ratings) and another modeling approach using a debiased machine learning approach (see Bajons, 2023).
@@
The remainder of the paper is structured as follows. In Section 2 we describe the data used for this research and the necessary preprocessing steps. In Section 3 we provide details on the novel regularization structures and models used to determine player strength. Section 4 presents a comparison of the ratings as obtained from the different models as well as use cases and an approach to statistically evaluate the rating methods. Finally, we conclude the paper with a discussion in Section 5.
@@ 2 Data
We use raw event stream data from the Spanish La Liga season of 2017/18, openly available via figshare (Pappalardo et al., 2019), as a basis in order to create a dataset of possessions. A possession is defined as a sequence of consecutive on-ball events, which ends either with the opponent team gaining the possession or by an action of the referee, i.e. a foul, the ball leaving the pitch, or the end of a period. As the main goal is the valuation of players, it is necessary to extract information on which player was involved in a possession sequence. Extracting information on offensive players involved in a possession is straightforward from the data, however, information on the opposing players is desirable as well. Since event stream data only provides information on the on-ball actions, it is difficult to extract meaningful opponent player information. One way to incorporate such information is to simply record all players on the field for a possession. Thus the regression matrix $X_{mod}=(X_{N_{inv}},X_{N_{of}})$ in this work contains $N=N_{inv}+N_{of}$ columns and $N_{p}$ rows. Each of the involvement columns $N_{inv}$ is a binary variable for a specific player, that is 1 whenever the player was involved in the possession and 0 otherwise. Each of the $N_{of}$ columns indicates whether a specific player was on the field for the possession team (indicated by 1) or on the field for the defensive team (indicated by $-1$) [^2] In this work $N_{inv}=N_{of}$ corresponds to the number of players in the data set.. The number of rows $N_{p}$ depends on the number of possessions analyzed. As we are interested in deriving a player rating it makes sense to only consider valuable possessions, where we define a possession as valuable if it ends in the last third of the pitch (i.e. sufficiently close to the opponent’s goal). Note that because we are interested in evaluating players we do not extract possession-specific variables. From the data, it would be possible to extract spatial and temporal features such as the length and speed of a possession. Thus, if we were solely interested in predicting the success of a possession it would probably be important to take such variables into account. However, the underlying assumption is that such spatial or temporal effects are due to the players being involved in the possession as well as on the field. That is, while it may be the case that quick but long possessions increase the chance of a successful outcome it is due to the quality of players involved and on the field that it is even possible to execute such an attack.
@@
Finally, a relevant response variable is extracted, which allows us to analyze the effect of players on possessions. Since goals are the most important part of the game of football, a goal indicator is taken as outcome variable. As football is a low-scoring game, one drawback is that this outcome variable is quite imbalanced. While there are other possibilities for outcome variables, such as ordinal type variables or hybrid approaches, it is beyond the scope of this paper to consider such cases.
The final data set used in this analysis covers possessions obtained from all of the 380 matches played in the 2017/18 season of the Spanish La Liga. Following our domain-specific intuition, only possessions that are comprised of more than 3 consecutive actions are considered. The only exceptions to this rule are possessions that start with a free kick or a corner kick, as they are relevant for scoring goals and often comprise only two actions (e.g. the corner kick and a shot). In most other circumstances, short possessions are a result of duels for the ball and are thus not necessarily directly relevant for detecting a player’s influence on scoring a goal from a possession. Discarding such possessions decreases the number to be analyzed from 141973 to $N_{p}=53233$ possessions. This, in particular, eliminates all the penalty kicks in our data [^3] It is common practice to separate goals scored from penalty kicks from goals scored within the flow of the game when analyzing the ability of football players, compare e.g. Robberechts and Davis, 2020 and Anzer and Bauer, 2021.. As we measure goals as outcome variable and these are sparse, a different definition of possessions or a different data truncation scheme may be considered in the future. For this work, we stick to the approach described, as this amounts to an increase in the percentage of possessions resulting in a goal from 0.7 % to 1.3 %. The final number of players in our data is 556 players, i.e. $N_{inv}=N_{of}=556$ and thus $N=1112$. On average the possessions contain $\approx 7.5$ consecutive actions with a median of $6$ consecutive actions. The average number of direct involvements for players in possession is 394.8. The distribution of the number of involvements is skewed to the right, which intuitively reflects the idea that only a few outstanding players have a lot of involvements whereas especially defenders are not as often involved in possessions. It is important to mention, that from a domain-specific viewpoint, several possible adjustments to the data setup could be considered. While we discuss some of these in more detail in Section 5, we emphasize that for the purpose of presenting the benefits of our novel extensions of top-down ratings, the setup is well-suited.
3 Methodology
@@
In the following, we discuss several approaches to deriving a rating of football players from the processed event data that overcome the previously mentioned problems. The main class of models is regularized or penalized regression, often referred to as adjusted plus-minus models, which have been studied in the context of sports player ratings by many authors previously (e.g., Gramacy et al., 2013; Hvattum, 2019; Kharrat et al., 2020). We adapt the methodology to our data setup and discuss various forms of penalized regression models. For each of the models, first, a statistical description is provided and secondly, a detailed domain-specific interpretation of the penalization structure as well as practical implications and use cases for practitioners are discussed.
@@ 3.1 Ridge regression
One of the earliest approaches for regularization dating back to Hoerl and Kennard (1970) is ridge regression (sometimes also referred to as Tikhonov regularization). The ridge estimator is given
@@
where $\ell(Y;X,\beta)$ represents the (negative) log-likelihood function for the binomial response variable $Y$ in the usual generalized linear model framework that we work in. In the case of a Gaussian likelihood function, an explicit solution for the above problem can be derived (see e.g. Hastie et al., 2013). For generalized linear models, the problem is typically formulated as a weighted least-squares problem which is then solved in an iterative fashion (cf. Friedman et al., 2010). The solution $\hat{\beta}_{ridge}$ of problem (1) is quite appealing as it shrinks the coefficients towards each other, where more shrinkage is put on the columns that explain less variance in the regression matrix $X$. That is, as we expect the response variable, i.e. goals scored from possession, to vary most in the directions of important players, this approach shrinks regular players more than strong (or weak) players. However, players with very similar strength will likely be shrunken towards a similar value. This is quite desirable from a perspective of a rating, as every player gets a rating that reflects his strengths in relation to every other player on the field.
Another particularly nice feature arises from the data setup. As the regression matrix $X=(X_{N_{inv}},X_{N_{of}})$ consists of involvement columns $N_{inv}$ of each player as well as of field indicator columns $N_{of}$, there are two sources of information on each player’s strengths. One is the direct information of how much a player $i$ contributes to specific possessions when involved, given by the corresponding (shrunken) estimate representing player $i$’s column in the sub-matrix $X_{N_{inv}}$. The other one is the indirect contribution to possessions, given by player $i$’s estimate from the sub-matrix $X_{N_{of}}$. From a practical perspective this is nicely interpretable: A player, even though potentially not involved in the possession, usually nevertheless has an influence on the game. A threatening forward for example may create space for other players, simply because the opposing team tries to cover the player. On the other hand, a strong goalkeeper who deflects a lot of shots also clearly influences offensive possessions of the opposing team. Such an effect may be measured from the estimate of indirect involvement of a player to a possession. The total strength of a player could then e.g. be measured by the sum of both contributions. In section 4 we discuss several methods of assembling the two strength estimates of a player in order to obtain an overall strength estimate of players.
@@ 3.2 Group lasso and exclusive lasso
A further very popular form of regularization is lasso regression, where instead of imposing an $L_{2}$-norm constraint on the coefficients $\beta$ (as is done in ridge regression), an $L_{1}$-norm penalty is used. The estimate is thus defined as
@@
This penalty allows for implicit variable selection, as some of the estimates are not only shrunken towards zero but set exactly to zero (cf. e.g. Hastie et al., 2013). While such an effect is desirable in many situations, as variables with less influence are omitted from the regression model and only influential variables retain a value different from zero, in our setup it is suboptimal. Even though it may be possible to detect strong (or weak) players from the resulting coefficients, the majority of them will be set to zero and no interpretation is possible anymore. This is especially inconvenient for the case of recruiting players, as a team is usually looking for a tradeoff between a strong player and an affordable player. Furthermore, the appealing effect of interpreting direct and indirect involvement to possession as in the ridge regression case is not as easy to employ. In the following, we discuss variants of the lasso penalty, which are more useful for our player evaluation framework.
@@
where the predictor variables are divided into $m$ groups. In equation (3), $\ell(Y;X,\beta)$ represents a scaled version of the negative log-likelihood and $p_{l}$ are weights representing the length of the $l$-th group specific estimate $\beta^{(l)}$ (cf. Simon et al., 2013). This penalty results in a shrinkage of the coefficients of complete groups to zero, i.e. on a group level it performs shrinkage similar to lasso. However, within each group, it does not perform variable selection, leading to a similar effect as ridge regression. Thus, if each group is of size one the group lasso results in the classical lasso. On the other hand, if there is only one group the penalty is similar to ridge regression.
In our application of rating football players, there is a natural grouping of players as usually, each player plays a specific position. While these positions may vary slightly depending on specific matches and tactics used, players can be categorized into at least 4 broad categories: Forward, midfielder, defender, and goalkeeper. The group lasso enables us to distinguish between players in each of these groups. On a group level, whole groups are shrunken simultaneously. If there were one irrelevant group (say for example the goalkeepers, whose main task is not to be involved in possession of high threat), then this group would be shrunken towards zero. Within a group however, the ridge-like penalty shrinks players of similar strength to a similar value, while less shrinkage is applied to strong (or weak) players. From an interpretational viewpoint, this is highly desirable. As the tasks on the field of each individual group vary, we do not expect comparisons between two groups, e.g. forwards and defenders, to be fair and instead might be more interested in within-group comparisons. Another dimension arises again from the distinction between direct and indirect involvement. Similar to the ridge regression case one can combine both aspects for the groups and perform within-group analyses.
@@
where $\ell(Y;X,\beta)$ again represents a scaled version of an exponential family likelihood function and the regressors can be separated into $m$ different groups. The behavior of this penalty is analyzed in detail in Campbell and Allen (2017). As opposed to the group lasso, it can be obtained from equation (4) that between groups an $L_{2}$ (i.e. ridge) penalty is applied, while within each group an $L_{1}$ (i.e. lasso) penalty is applied. That is, on a group level the penalty does not imply shrinkage of complete groups, i.e. in each group at least one variable will be non-zero. Within groups however only the most relevant variables will be selected, i.e. the penalty tries to set most of the group members to zero as is usual for the lasso penalty. In the extreme case, exactly one non-zero variable of each group will be selected.
The application-specific use case of the penalty reveals more potential in scouting rather than in rating players. While one could again categorize the players into the above 4 categories based on their broad position on the pitch and analyze which players are particularly strong (resp. weak) in their respective group, the exclusive lasso motivates more interesting selection procedures. A particular use case would be a team that is interested in signing players who play a specific position and who are at a certain age level or below a certain market value. In this case, an individual group could be formed and the exclusive lasso would result in the selection of a few strong (resp. weak) players that fall into the prespecified type, while at the same time being able to account for the usual groups as defined by e.g. player position. More details on using the exclusive lasso for scouting are found in Appendix B for this work.

### v2

@@ Abstract
This paper presents a novel framework for evaluating players in association football (soccer). Our method uses possession sequences, i.e.~sequences of consecutive on-ball actions, for deriving estimates for player strengths. On the surface, the methodology is similar to classical adjusted plus-minus rating models using mainly regularized regression techniques. However, by analyzing possessions, our framework is able to distinguish on-ball and off-ball contributions of players to the game. From a methodological viewpoint, the framework explores four different penalization schemes, which exploit football-specific structures such as the grouping of players into position groups as well as into common strength groups. These four models lead to four ways to rate players by considering the respective estimate of each model corresponding to the player. The ratings are used to analyze the 2017/18 season of the Spanish La Liga. We compare similarities as well as particular use cases of each of the penalized models and provide guidance for practitioners when using the individual model specifications. Finally, we conclude our analysis by providing a domain-specific statistical evaluation framework, which highlights the potential of the penalized regression approaches for evaluating players.
@@ 1 Introduction
In football, the decision-making process is traditionally realized by a qualitative assessment of players by professionals such as coaches, managers, or scouts. However, lately, there have been many attempts and advances toward a more objective approach to quantifying the impact of players on the game. For example, traditional statistics such as goals or pass completion percentage, which have been used extensively to analyze the strength of players, have been replaced by metrics such as expected goals (xG, cf.e.g.Robberechts and Davis, 2020; Anzer and Bauer, 2021) or passing models based on statistical approaches (Szczepański and McHale, 2016; Håland et al., 2019, 2020). In general, these novel data-driven methods can be categorized into two distinct groups: bottom-up and top-down rating systems (cf.~Hvattum and Gelade, 2021). The former assesses player performances by assigning values to each action performed and then aggregating them for each player over the course of a relevant period (i.e.~a match or a season). Top-down ratings on the other hand evaluate players by breaking down the whole team’s performance and distributing credit onto the players involved. Typically, bottom-up approaches use very granular data on an event level in order to assign values to either specific actions such as shots (xG) or passes or more generally any type of action (cf.~Decroos et al., 2019). While bottom-up ratings are therefore able to assess a player’s quality down to each particular action performed by the athlete, they however lack the ability to account for team strengths as well as off-ball actions by teammates. In contrast, top-down ratings such as plus-minus (PM) models (see Hvattum, 2019 for an overview) rely only on information gained from specific segments of a match, which in general is much less detailed than event stream data. These approaches are by design not able to account for specific actions taken by a player. They are, however, able to incorporate effects of teammates.
@@
In this work, we present a novel framework for evaluating player strength. On the surface, the methodology is similar to regularized adjusted plus-minus (PM) models as we mainly use penalized regression models in order to infer a rating of players. However, we extend the classical PM approach in two ways. First, we separate matches into possessions, which are sequences of consecutive on-ball actions. Thus, we consider much more granular segments of the game than existing PM models, where a segment is defined as one where the same set of players is on the field (Kharrat et al., 2020; Hvattum, 2019). The segmentation in possessions leads to a data structure where it is possible to separate effects of being part of a possession, i.e.~actually performing on-ball actions (direct involvement), from effects of only being on the pitch (indirect involvement). From a domain-specific viewpoint, this is highly attractive as it is well-known that off-ball actions have a substantial impact in soccer (cf. Spearman, 2018; Fernández et al., 2021). Using possessions data, it is also necessary to adjust the dependent variable that measures the outcome of a segment. In the classical regularized adjusted PM approach the (normalized) goal difference in a segment is used. Since possessions are way shorter segments, goal difference is not appropriate as response variable. Instead, we focus on whether the possession ends in a goal or not as our outcome variable. Second, we provide a variety of penalization structures, which have not been considered in this context. Most of the classical plus-minus models use a ridge regression penalty, i.e.~an $L_{2}$-type of penalty on the regression coefficient (cf.~e.g. Kharrat et al., 2020). In a more innovative approach, Pantuso and Hvattum (2021) adjust the penalty term by accounting for an age effect as well as a similarity effect within the shrinkage procedure. However, to the best of our knowledge, there has been no comparison or analysis of penalization structures beyond (adaptions of) the ridge penalty. We analyze several regularization structures that exploit the implicit grouping of players that arises from the different position groups on the field. In this way, our player evaluations are able to provide interesting and novel insights into player strengths where different penalties may be used for different use cases from practitioners. Finally, we conduct a statistical evaluation of our methods specifically suited for football based on validity experiments from Hvattum and Gelade (2021). Our methods are compared to reference models using different kinds of data (ELO team ratings) and another modeling approach using a debiased machine learning approach (Bajons, 2023).
@@
The remainder of the paper is structured as follows. In Section 2 we describe the data used for this research and the necessary preprocessing steps. In Section 3, we provide details on the novel regularization structures and models used to determine player strength. Statistical descriptions are accompanied by managerial implications of each model. Section 4 presents a comparison of the ratings as obtained from the different models and our approach to statistically evaluate the methods. Finally, we conclude the paper with a discussion in Section 5.
@@ 2 Data
We use raw event stream data from the Spanish La Liga season of 2017/18, openly available via Figshare (Pappalardo et al., 2019), as a basis in order to create a dataset of possessions. A possession is defined as a sequence of consecutive on-ball events, which ends either with the opponent team gaining the possession or by an action of the referee, i.e.~a foul, the ball leaving the pitch, or the end of a period.
@@
As the main goal is the valuation of players, it is necessary to extract information on which player was involved in a possession sequence. Extracting offensive player involvement in possessions is straightforward from the data, however, information on the opposing players is desirable as well. Since event stream data only provides details on the on-ball actions, it is difficult to extract meaningful opponent player information. One way to incorporate such knowledge is to record all players on the field during a possession. Thus, we consider a regression matrix $X_{mod}=(X_{N_{inv}},X_{N_{of}})$ with $N=N_{inv}+N_{of}$ columns and $N_{p}$ rows. Each of the involvement columns $N_{inv}$ is a binary variable for a specific player, that is 1 whenever the player was involved in the possession and 0 otherwise. Each of the $N_{of}$ columns indicates whether a specific player was on the field for the possession team (indicated by 1) or on the field for the defensive team (indicated by $-1$). In this work, $N_{inv}=N_{of}=556$ corresponds to the number of players in the data set. Note that because we are interested in evaluating players we do not extract possession-specific variables. The data includes spatial and temporal features such as the length and speed of possessions, which may help predict the success of these. However, the underlying assumption is that spatial or temporal effects are due to the players being involved in the possession as well as on the field. That is, while it may be the case that quick but long possessions increase the chance of a successful outcome it is due to the quality of players involved and on the field that it is even possible to execute such an attack.
The number of rows $N_{p}$ depends on the number of possessions analyzed. As we are interested in deriving a player rating it makes sense to only consider valuable possessions, where we define a possession as valuable if it ends in the last third of the pitch (i.e.~sufficiently close to the opponent’s goal). Following our domain-specific intuition, we further only consider possessions with at least 3 consecutive actions. The only exceptions to this rule are possessions that start with a free kick or a corner kick, as they are relevant for scoring goals and often comprise only two actions (e.g.~the corner kick and a header). This, in particular, eliminates all the penalty kicks in our data. It is common practice to exclude these highly specific situations (Robberechts and Davis, 2020; Anzer and Bauer, 2021). In most other circumstances, short possessions are a result of duels for the ball and are thus not necessarily directly relevant for detecting a player’s influence on scoring a goal from a possession. In total, the data contains $N_{p}=53233$ possessions.
@@
Finally, a relevant response variable is extracted, which allows us to analyze the effect of players on possessions. Since goals are the most important part of the game of football, a goal indicator is taken as outcome variable. As football is a low-scoring game, one drawback is that this outcome variable is quite imbalanced. In our data, only 1.3 % of the possessions end in a goal.
On average the possessions contain $\approx 7.5$ consecutive actions with a median of $6$ consecutive actions. The average number of direct involvements for players in possession is 394.8. The distribution of the number of involvements is skewed to the right, which intuitively reflects the idea that only a few outstanding players have a lot of involvements whereas especially defenders are not as often involved in possessions. It is important to mention, that from a domain-specific viewpoint, several possible adjustments to the data setup could be considered. While we discuss some of these in more detail in Section 5, we emphasize that for the purpose of presenting the benefits of our methods, the setup is well-suited.
3 Methods and practical implications
@@
In the following, we discuss several approaches to deriving a rating of football players from the processed event data that overcome the previously mentioned problems. The main class of models is regularized or penalized regression, often referred to as adjusted plus-minus models, which have been studied in the context of sports player ratings by many authors previously (e.g., Gramacy et al., 2013; Hvattum, 2019; Kharrat et al., 2020). We adapt the methodology to our data setup and discuss various forms of penalized regression models. For each of the models, first, a statistical description is provided and secondly, a detailed domain-specific interpretation of the penalization structure as well as implications and use cases for practitioners are discussed.
@@ 3.1 Ridge regression
One of the earliest approaches for regularization dating back to Hoerl and Kennard (1970) is ridge regression (sometimes also referred to as Tikhonov regularization). The ridge estimator is given as
@@
where $\ell(Y;X,\beta)$ represents the (negative) log-likelihood function for the binomial response variable $Y$ in the usual generalized linear model framework that we work in. % In the case of a Gaussian likelihood function, an explicit solution for the above problem can be derived (see e.g. Hastie et al., 2013). For generalized linear models, the problem is typically formulated as a weighted least-squares problem which is then solved in an iterative fashion (cf. Friedman et al., 2010). The solution $\hat{\beta}_{ridge}$ of problem (1) is quite appealing as it shrinks the coefficients towards each other, where more shrinkage is put on the columns that explain less variance in the regression matrix $X$. That is, as we expect the response variable, i.e.~goals scored from possession, to vary most in the directions of important players, this approach shrinks regular players more than strong (or weak) players. However, players with very similar strength will likely be shrunken towards a similar value. This is quite desirable from a perspective of a rating, as every player gets a rating that reflects his strengths in relation to every other player on the field.
Another particularly nice feature arises from the data setup. As the regression matrix $X=(X_{N_{inv}},X_{N_{of}})$ consists of direct involvement columns $N_{inv}$ as well as on-field indicator columns $N_{of}$, there are two sources of information for each player $i$ represented by two estimates $\hat{\beta}_{i,inv}$ and $\hat{\beta}_{i,of}$. %The (shrunken) estimate $\hat{\beta}_{i,inv}$ of player $i$ represents a direct involvement strenght, while $\hat{\beta}_{i,of}$ characterized the effect of player $i$ being on the field. %One is the direct information of how much a player $i$ contributes to specific possessions when involved, given by the corresponding (shrunken) estimate representing player $i$’s column in the sub-matrix $X_{N_{inv}}$. The other one is the indirect contribution to possessions, given by player $i$’s estimate from the sub-matrix $X_{N_{of}}$. From a practical perspective this is nicely interpretable: A player, even though potentially not involved in the possession, usually nevertheless has an influence on the game. A threatening forward for example may create space for other players, simply because the opposing team tries to cover the player. On the other hand, a strong goalkeeper who deflects a lot of shots also clearly influences offensive possessions of the opposing team. Such effects may be measured from the estimate of indirect involvement $\hat{\beta}_{i,of}$. The total strength of a player $i$ could then e.g.~be measured by the sum of both contributions. In Section 4 we discuss several methods of assembling the two strength estimates of a player in order to obtain an overall strength estimate of players.
@@ 3.2 Group lasso and exclusive lasso
Another common form of regularization is lasso regression, where instead of imposing an $L_{2}$-norm constraint on the coefficients $\beta$ (as is done in ridge regression), an $L_{1}$-norm penalty is used. The estimate is thus defined as
@@
This penalty allows for implicit variable selection, as some of the estimates are not only shrunken towards zero but set exactly to zero (cf.~e.g. Hastie et al., 2013). While such an effect is desirable in many situations, as variables with less influence are omitted from the regression model and only influential variables retain a value different from zero, in our setup it is suboptimal. Even though it may be possible to detect strong (or weak) players from the resulting coefficients, the majority of them will be set to zero and no interpretation is possible anymore. This is especially inconvenient for the case of recruiting players, as a team is usually looking for a tradeoff between a strong player and an affordable player. Furthermore, the appealing effect of interpreting direct and indirect involvement to possession as in the ridge regression case is not as easy to employ. In the following, we discuss variants of the lasso penalty, which are more useful for our player evaluation framework.
@@
where the predictor variables are divided into $m$ groups. In equation (3), $\ell(Y;X,\beta)$ represents a scaled version of the negative log-likelihood and $p_{l}$ are weights representing the length of the $l$-th group specific estimate $\beta^{(l)}$ (cf. Simon et al., 2013). This penalty results in a shrinkage of the coefficients of complete groups to zero, i.e.~on a group level it performs shrinkage similar to lasso. However, within each group, it does not perform variable selection, leading to a similar effect as ridge regression. Thus, if each group is of size one the group lasso results in the classical lasso. On the other hand, if there is only one group the penalty is similar to ridge regression.
In our application of rating football players, there is a natural grouping of players as usually, each player plays a specific position. While these positions may vary slightly depending on specific matches and tactics used, players can be categorized into at least 4 broad categories: Forward, midfielder, defender, and goalkeeper. The group lasso enables us to distinguish between players in each of these groups. On a group level, whole groups are shrunken simultaneously. If there were one irrelevant group (say for example the goalkeepers, whose main task is not to be involved in possession of high threat), then this group would be shrunken towards zero. Within a group however, the ridge-like penalty shrinks players of similar strength to a similar value, while less shrinkage is applied to strong (or weak) players. From an interpretational viewpoint, this is highly desirable. As the tasks on the field of each individual group vary, we do not expect comparisons between two groups, e.g.~forwards and defenders, to be fair and instead might be more interested in within-group comparisons. Another dimension arises again from the distinction between direct and indirect involvement. Similar to the ridge regression case one can combine both aspects for the groups and perform within-group analyses.
@@
where $\ell(Y;X,\beta)$ again represents a scaled version of an exponential family likelihood function and the regressors can be separated into $m$ different groups. The behavior of this penalty is analyzed in detail in Campbell and Allen (2017). As opposed to the group lasso, it can be obtained from equation (4) that between groups an $L_{2}$ (i.e.~ridge) penalty is applied, while within each group an $L_{1}$ (i.e.~lasso) penalty is applied. That is, on a group level the penalty does not imply shrinkage of complete groups, i.e.~in each group at least one variable will be non-zero. Within groups however only the most relevant variables will be selected, i.e.~the penalty tries to set most of the group members to zero as is usual for the lasso penalty. In the extreme case, exactly one non-zero variable of each group will be selected.
The application-specific use case of the penalty reveals more potential in scouting rather than in rating players. While one could again categorize the players into the above 4 categories based on their broad position on the pitch and analyze which players are particularly strong (resp. weak) in their respective group, the exclusive lasso motivates more interesting selection procedures. A particular use case would be a team that is interested in signing players who play a specific position and who are at a certain age level or below a certain market value. In this case, an individual group could be formed and the exclusive lasso would result in the selection of a few strong (resp. weak) players that fall into the prespecified type, while at the same time being able to account for the usual groups as defined by e.g.~player position. More details on using the exclusive lasso for scouting are found in the supplementary materials for this work.